OpenStudy (deadshot):

Simplify quantity 12 x minus 96 over quantity x squared minus 4x minus 32 all over quantity 6 x squared minus 6 x minus 12 over 3 x plus 3

OpenStudy (anonymous):

it's C

OpenStudy (deadshot):

so it's C. open parentheses x plus 4 close parentheses open parentheses x minus 2 close parentheses over 6?

OpenStudy (anonymous):

can you write this with the equation tool?

OpenStudy (deadshot):

The question is\[\frac{ \frac{ 12x-96 }{ x^2-4x-32 } }{ \frac{ 6x^2-6x-12 }{ 3x+3 } }\] and answer c is \[\frac{ (x+4)(x-2) }{ 6 }\]